# Beatty solutions of almost Golomb equations

### 16 Apr 2026 v2

@@ 4.1 The continuous perspective
4.2 Expansion of the triple composition
Throughout this section, $c=1/\!\sqrt{2}$, $d\in\mathbb{R}$, and $a(n)=\lfloor cn+d\rfloor$. Write $e(n)=-\{cn+d\}\in(-1,0]$, so that $a(n)=cn+d+e(n)$. Define f(n)=a(n)+a(n1)f(n)=a(n)+a(n{-}1). Then
@@ 4.3 Necessity
Any solution must satisfy $a(1)=\lfloor c+d\rfloor=1$, requiring $1-c\leq d<2-c$. The failure of (5) at both extremes of $I$ is detected along the continued-fraction convergents of $c$. We record the needed Pell data.
Lemma 15.
Let $p_{k}/q_{k}$ be the $k$-th convergent of $c=[0;1,2,2,2,\ldots]$ and set $\varepsilon_{k}=(-1)^{k}(cq_{k}-p_{k})>0$. The first values are
$k$ 11 22 33 44 55 66 77 88
$p_{k}$ 11 22 55 $12$ $29$ $70$ $169$ $408$
$q_{k}$ 11 33 77 $17$ $41$ $99$ $239$ $577$
$\varepsilon_{k}$ .293.293 .121.121 .050.050 .021.021 .009.009 .004.004 .001.001 .0006.0006
and the following identities hold for every $k\geq 1$:
  1. 1.
    $cq_{k}=p_{k}+(-1)^{k}\varepsilon_{k}$ with $\varepsilon_{k}\to 0^{+}$;
  2. 2.
    $2cp_{k}=q_{k}-2c(-1)^{k}\varepsilon_{k}$.
Proof.
The numerical table is standard. Identity (i) is the definition of $\varepsilon_{k}$. Identity (ii) follows from (i): $2cp_{k}=2c(cq_{k}-(-1)^{k}\varepsilon_{k})=2c^{2}q_{k}-2c(-1)^{k}\varepsilon_{k}=q_{k}-2c(-1)^{k}\varepsilon_{k}$, using $2c^{2}=1$. ∎
@@
Write $d=c-\delta$ with $\delta>0$. Fix $k\geq 2$ even large enough that $\varepsilon_{k}<\delta$, so that (i) gives $cq_{k}=p_{k}+\varepsilon_{k}$ with $0<\varepsilon_{k}<\delta<c<1$. Evaluate at $n=q_{k}$:
$a(q_{k})$ $=\lfloor cq_{k}+d\rfloor=\lfloor p_{k}+\varepsilon_{k}+c-\delta\rfloor=p_{k},$
$a(q_{k}{-}1)$ $=\lfloor c(q_{k}{-}1)+d\rfloor=\lfloor p_{k}+\varepsilon_{k}-\delta\rfloor=p_{k}-1,$
where the first floor uses $\varepsilon_{k}+c-\delta\in(0,c)\subset(0,1)$ and the second uses $1<\varepsilon_{k}-\delta<0$. Hence $f(q_{k})=a(q_{k})+a(q_{k}{-}1)=2p_{k}-1$, and by identity (ii):
$a(f(q_{k}))=\lfloor c(2p_{k}-1)+d\rfloor=\lfloor q_{k}-2c\varepsilon_{k}-\delta\rfloor=q_{k}-1,$
since $0<2c\varepsilon_{k}+\delta<1$ for $k$ large. Applying $a$ once more:
$a(a(f(q_{k})))=a(q_{k}-1)=p_{k}-1\neq p_{k}=a(q_{k}),$
which violates (5) at $n=q_{k}$. Since infinitely many even $k$ are admissible, (5) fails at infinitely many $n$. ∎
@@
Write $u=4c-2=2(\sqrt{2}-1)$ and $d=u+\delta$ with $\delta>0$. The bound d<2cd<2-c forces $\delta<1-c$. Fix $k\geq 3$ odd large enough that $\varepsilon_{k}<\delta$ and $2c\varepsilon_{k}+\delta<1-c$, and take $n=q_{k}-3$. Identity (i) with $k$ odd gives $cq_{k}=p_{k}-\varepsilon_{k}$, hence
$a(q_{k}{-}3)$ $=\lfloor c(q_{k}{-}3)+d\rfloor=\lfloor p_{k}-\varepsilon_{k}-3c+u+\delta\rfloor$
$=\lfloor p_{k}-(2-c)+(\delta-\varepsilon_{k})\rfloor=p_{k}-2,$
since $2-c\approx 1.293$ and $0<\delta-\varepsilon_{k}<1-c$, so $(2-c)+(\delta-\varepsilon_{k})$ lies in $(c-2,c-1)\subset(-2,-1)$, giving floor value $p_{k}-2$. Similarly
$a(q_{k}{-}4)=\lfloor p_{k}-(2-2c)+(\delta-\varepsilon_{k})\rfloor=p_{k}-2,$
since $(2-2c)+(\delta-\varepsilon_{k})$ lies in $(2c-2,2c-1)\subset(-1,0)$. Hence $q_{k}-3$ is a repeat and $f(q_{k}{-}3)=2p_{k}-4$. By identity (ii),
$a(2p_{k}{-}4)$ $=\lfloor c(2p_{k}{-}4)+d\rfloor=\lfloor q_{k}+2c\varepsilon_{k}-4c+u+\delta\rfloor$
$=\lfloor q_{k}-(2-2c\varepsilon_{k}-\delta)\rfloor=q_{k}-2,$
since $2c\varepsilon_{k}+\delta<1-c<1$, so $(2-2c\varepsilon_{k}-\delta)\in(-2,-1)$. Finally
$a(q_{k}{-}2)=\lfloor p_{k}-\varepsilon_{k}-2c+u+\delta\rfloor=\lfloor p_{k}+(2c-2)+(\delta-\varepsilon_{k})\rfloor=p_{k}-1,$
since $(2c-2)+(\delta-\varepsilon_{k})\in(2c-2,1-2c)\subset(-1,0)$. Therefore
$a(a(f(q_{k}{-}3)))=a(q_{k}{-}2)=p_{k}-1\neq p_{k}-2=a(q_{k}{-}3),$
which violates (5) at $n=q_{k}-3$. Infinitely many odd $k$ are admissible. ∎
@@ 4.4 Sufficiency
Proof of Theorem 3.
Write $\theta=\{cn+d\}$, so $a(n)=cn+d-\theta$. The index $n$ is a repeat if $\theta<1{-}c$ (giving a(n+1)=a(n)a(n{+}1)=a(n)) and a non-repeat if $\theta\geq 1{-}c$. Set D=d(2c+1)D=d(2c{+}1).
@@
Non-repeat, $\theta\in[c,1)$. Then $f(n)=2a(n)$ and $a(f(n))=\lfloor n+\varphi\rfloor$ with $\varphi=D-2c\theta\in(D-2c,\;D-1]$. Since $d\geq c$ gives $\varphi>0$ and $d\leq 4c{-}2$ gives $\varphi\leq 1$ with equality only when d=4c2d=4c{-}2 and $\theta=c$ (requiring $c(n{+}3)\in\mathbb{Z}$, impossible), $\varphi\in(0,1)$ and $a(f(n))=n$.
@@
Non-repeat, $\theta\in[1{-}c,c)$. Then f(n)=2a(n)1f(n)=2a(n)-1 and $\varphi=D-c(2\theta{+}1)\in(0,1)$ by the same bounds on $D$.
@@
Repeat ($\theta\in[0,1{-}c)$). Then $\varphi\in(0,2)$, so $a(f(n))\in\{n,n{+}1\}$. Since $n$ is a repeat, a(n+1)=a(n)a(n{+}1)=a(n), hence $a(a(f(n)))=a(n)$. ∎
4.5 Wythoff endpoints
The endpoints of the Beatty interval I=[2/2, 2(21)]I=[$\sqrt${2}/2,\;2($\sqrt${2}{-}1)] coincide with the parameters of the Wythoff complementary pair.

### 21 Apr 2026 v3

@@ 4.1 The continuous equation
4.2 Expansion
Throughout this section, $c=1/\!\sqrt{2}$, $d\in\mathbb{R}$, and $a(n)=\lfloor cn+d\rfloor$. Write $\theta_{n}=\{cn+d\}\in[0,1)$, so that $a(n)=cn+d-\theta_{n}$, and say that index $n$ is a repeat when a(n+1)=a(n)a(n{+}1)=a(n), and a non-repeat when a(n+1)=a(n)+1a(n{+}1)=a(n){+}1. Since $a(n{+}1)-a(n)=\lfloor\theta_{n}+c\rfloor$, this amounts to $\theta_{n}<1{-}c$ (repeat) or $\theta_{n}\geq 1{-}c$ (non-repeat). Set also $e(n)=-\theta_{n}\in(-1,0]$, so that $a(n)=cn+d+e(n)$. Define f(n)=a(n)+a(n1)f(n)=a(n)+a(n{-}1). Then
@@ 4.3 Necessity
Any solution must satisfy $a(1)=\lfloor c+d\rfloor=1$, requiring $1-c\leq d<2-c$. The failure of (5) outside $I$ is produced by a local argument on the three residue classes of $\theta_{n}\in[0,1)$ combined with density of the orbit $(\theta_{n})_{n\geq 1}$. We describe the dynamics once, and then read off each endpoint as a special case.
Expansion (16) together with $2c^{2}=1$ gives
$a(f(n))=\lfloor n-\tfrac{1}{2}+D-c\Theta_{2}(n)\rfloor,\qquad D=d(2c+1),\quad\Theta_{2}(n)=\theta_{n}+\theta_{n-1}.$ (19)
Writing $a(f(n))=n+m(n)$, the integer $m(n)$ is determined by
$m(n)=k\iff D-c\Theta_{2}(n)\in\Bigl[k+\tfrac{1}{2},\,k+\tfrac{3}{2}\Bigr).$
Two simple implications are immediate. If $m(n)\geq 1$, then $a(f(n))\geq n+1$, so survival of the identity $a(a(f(n)))=a(n)$ forces $a(n+1)=a(n)$, that is, $n$ is a repeat. Symmetrically, if $m(n)\leq-1$ then n1n-1 must be a repeat. These implications are the two basic necessary conditions
  • (N+)
    $m(n)\geq 1\ \Rightarrow\ a(n+1)=a(n)$ ($n$ is a repeat),
  • (N-)
    $m(n)\leq-1\ \Rightarrow\ a(n-1)=a(n)$ (n1n-1 is a repeat).
Taken as a global statement over all $d$, (N+) and (N-) are only necessary. If $|m(n)|\geq 2$, a single repeat is no longer enough to absorb the extra jump, and the identity fails even when $n$ or n1n-1 happens to be a repeat. The bound $|m(n)|\leq 1$ must therefore be ensured a priori. We show in Section 4.4 that this bound holds uniformly on the interval II, so that on $I$ the two conditions (N+) and (N-) become jointly sufficient as well. The necessity lemmas below exploit the two implications directly.
The three residue classes of $\theta=\theta_{n}$ are
$\mathrm{I}:\ \theta\in[0,1{-}c),\qquad\mathrm{II}:\ \theta\in[1{-}c,c),\qquad\mathrm{III}:\ \theta\in[c,1),$
which match the three branches of the decomposition in Lemma 20 below. In class I, $\theta_{n}<1-c$ means $n$ is a repeat. In classes II and III, $n$ is a non-repeat, so (N+) forbids $m(n)\geq 1$ there. In class II, furthermore, $\theta_{n-1}=\theta-c+1\in[2-2c,1)\subset[1-c,1)$, so n1n-1 is also a non-repeat, and (N-) forbids $m(n)\leq-1$ on class II.
Write $\Theta_{2}(n)$ explicitly in each class. In class III, $\theta_{n-1}=\theta-c\in[0,1-c)$, hence $\Theta_{2}(n)=2\theta-c$ and
$c\Theta_{2}(n)=2c\theta-\tfrac{1}{2}.$ (20)
In class II, $\theta_{n-1}=\theta+1-c$, hence $\Theta_{2}(n)=2\theta+1-c$ and
$c\Theta_{2}(n)=2c\theta+c-\tfrac{1}{2}.$ (21)
The orbit $(\theta_{n})_{n\geq 1}$ is the forward orbit of $\{c+d\}$ under the irrational rotation $\theta\mapsto\{\theta+c\}$. By Weyl’s equidistribution theorem, this orbit is equidistributed in $[0,1)$, so it visits every non-empty open sub-interval of $[0,1)$ at infinitely many indices $n$.
@@
Compute $D$ at $d=c$: $D=c(2c+1)=2c^{2}+c=1+c$. So $d<c$ is equivalent to $D<1+c$. Work in class II, where $\theta\in[1{-}c,c)$ and $c\Theta_{2}=2c\theta+c-1/2$ by (21). As $\theta\to c^{-}$, $c\Theta_{2}\to 2c^{2}+c-1/2=c+1/2$, and so
$D-c\Theta_{2}\ \longrightarrow\ D-c-\tfrac{1}{2}.$
The hypothesis $D<1+c$ gives Dc1/2<1/2D-c-1/2<1/2, so the inequality $D-c\Theta_{2}(n)<1/2$ defines an open sub-interval of class II accumulating at $\theta=c$. Explicitly, solving $D-c\Theta_{2}<1/2$ for $\theta$ yields
$\theta>\frac{D-c}{2c},$ (22)
and the right-hand side is <c<c by hypothesis. For every $n$ with $\theta_{n}$ in this sub-interval, $m(n)\leq-1$ (and in fact m(n)=1m(n)=-1, since $D\geq c$ from the lower bound $d\geq 1-c$ keeps $D-c\Theta_{2}$ above 1/2-1/2 in this range).
Both $n$ and n1n-1 are non-repeats throughout class II, so (N-) is violated at each such $n$. Density of the orbit provides infinitely many such $n$, and (5) fails at infinitely many $n$. ∎
@@
Set $d_{R}=2(\sqrt{2}-1)=4c-2$. Compute $D$ at $d_{R}$:
$D(d_{R})=(4c-2)(2c+1)=8c^{2}+4c-4c-2=4-2=2,$
using $2c^{2}=1$. So $d>d_{R}$ is equivalent to $D>2$. Work in class III, where $\theta\in[c,1)$ and $c\Theta_{2}=2c\theta-1/2$ by (20). As $\theta\to c^{+}$, $c\Theta_{2}\to 2c^{2}-1/2=1/2$, and so
$D-c\Theta_{2}\ \longrightarrow\ D-\tfrac{1}{2}.$
The hypothesis $D>2$ gives D1/2>3/2D-1/2>3/2, so the inequality $D-c\Theta_{2}(n)\geq 3/2$ defines an open sub-interval of class III accumulating at $\theta=c$. Explicitly, solving $D-c\Theta_{2}\geq 3/2$ for $\theta$ yields
$\theta\leq\frac{D-1}{2c},$ (23)
and the right-hand side is >c>c by hypothesis. For every $n$ with $\theta_{n}$ in this sub-interval, $m(n)\geq 1$.
Throughout class III, $n$ is a non-repeat, so (N+) is violated. Density of the orbit supplies infinitely many such $n$, and (5) fails at infinitely many $n$. ∎
Remark 19.
The two endpoints $d=c$ and $d=d_{R}$ survive because the limiting condition would require $\theta=c$ exactly. This forces $\{cn+d\}=c$ for some integer $n$, hence $c(n-1)+d\in\mathbb{Z}$, impossible since $c$ is irrational and $d\in\mathbb{Q}(\sqrt{2})\setminus\mathbb{Q}$ at both endpoints. The orbit is dense, and irrationality excludes the single critical address at each endpoint. This dynamical argument makes the two proofs symmetric and self-contained, despite the asymmetry of $I$.
@@ 4.4 Sufficiency
The sufficiency of $d\in I$ rests on an absorption mechanism: the first nesting $a(f(n))$ may exceed $n$ by one at repeat positions, and this excess is absorbed by the outer $a$ because $a(n+1)=a(n)$ at such positions. We isolate this mechanism as a lemma and then deduce the theorem.
Lemma 20.
Let $c=1/\!\sqrt{2}$, $a(n)=\lfloor cn+d\rfloor$, and f(n)=a(n)+a(n1)f(n)=a(n)+a(n-1). For every $d\in I$ and every $n\geq 2$, $a(f(n))\in\{n,n+1\}$, with the finer dichotomy
  • if $\theta_{n}\geq 1-c$ (i.e. $n$ is a non-repeat), then $a(f(n))=n$;
  • if $\theta_{n}<1-c$ (i.e. $n$ is a repeat), then $a(f(n))\in\{n,n+1\}$.
Proof.
Write $\theta=\theta_{n}$ and D=d(2c+1)D=d(2c{+}1). We distinguish three residue classes of $\theta$.
@@
Suppose first that $\theta\in[c,1)$. Then $f(n)=2a(n)$ and $a(f(n))=\lfloor n+\varphi\rfloor$ with $\varphi=D-2c\theta\in(D-2c,\;D-1]$. Since $d\geq c$ gives $\varphi>0$ and $d\leq 4c{-}2$ gives $\varphi\leq 1$ with equality only when d=4c2d=4c{-}2 and $\theta=c$ (requiring $c(n{+}3)\in\mathbb{Z}$, impossible), $\varphi\in(0,1)$ and $a(f(n))=n$.
@@
Suppose next that $\theta\in[1{-}c,c)$. Then f(n)=2a(n)1f(n)=2a(n)-1 and $\varphi=D-c(2\theta{+}1)\in(0,1)$ by the same bounds on $D$, hence $a(f(n))=n$. The two sub-cases cover non-repeat positions.
@@
Suppose finally that $\theta\in[0,1{-}c)$, so $n$ is a repeat. Then $\varphi\in(0,2)$, whence $a(f(n))\in\{n,n{+}1\}$. ∎
Proof of Theorems 3 and 16, sufficiency.
Lemma 20 gives $a(f(n))\in\{n,n+1\}$ for every $d\in I$ and $n\geq 2$, so the shift $m(n)$ introduced in Section 4.3 takes values in $\{0,1\}$. The a priori bound $|m(n)|\leq 1$ discussed there therefore holds on $I$, and the conditions (N+) and (N-) become sufficient as well as necessary. Concretely, the lemma gives $a(f(n))=n$ at non-repeat positions, so $a(a(f(n)))=a(n)$ trivially. At repeat positions $a(f(n))\in\{n,n+1\}$, and since $n$ is a repeat, a(n+1)=a(n)a(n{+}1)=a(n), so $a(a(f(n)))=a(n)$ in both subcases. The triple-nested identity therefore holds for every $n\geq 2$ and every $d\in I$. ∎
4.5 The Ostrowski–Pell framework behind the $r=2$ triple-nested family
The endpoint formulas, the return-time structure of the defect set, and the underlying irrational rotation are naturally expressed in the Ostrowski–Pell framework attached to the quadratic irrational $\alpha=1+\sqrt{2}$, organised by Fokkink [4] as a continuation of the Wythoff array of Conway and Ryba.
Shared continued-fraction backbone
The slope $c=1/\!\sqrt{2}$ has continued-fraction expansion
$c=[0;1,2,2,2,\ldots],$
whose convergents $p_{k}/q_{k}$ have denominators
$q_{k}=1,3,7,17,41,99,239,577,\ldots$
(A001333), the Pell denominators satisfying the recurrence $q_{k+1}=2q_{k}+q_{k-1}$. The characteristic equation $x^{2}=2x+1$ of this recurrence has dominant root $\alpha=1+\sqrt{2}$, the silver ratio. Fokkink’s Pell tower is built from the same recurrence as the $m=2$ instance of the family of bi-infinite sequences
$X_{n+1}=m\,X_{n}+X_{n-1},\qquad m\geq 1,$
where $m$ is a positive-integer parameter of the tower, unrelated to the shift $d$ of our Beatty ansatz. The case $m=1$ recovers the Fibonacci recurrence and the Wythoff array of Morrison and Kimberling.
Shared numeration backbone and the wall column
Every positive integer admits a unique Ostrowski representation
$n=\sum_{k\geq 1}e_{k}q_{k}$
with digits $e_{1}\in\{0,1\}$, $e_{k}\in\{0,1,2\}$ for $k\geq 2$, and $e_{k}=2\Rightarrow e_{k-1}=0$. In the tower picture, this numeration labels the rows. Fokkink distinguishes two natural Beatty sequences attached to the tower. The first column $A_{1,n}$ is the non-homogeneous Beatty sequence
$A_{1,n}=\Bigl\lfloor\frac{n\alpha}{\alpha-1}-\frac{1}{\alpha(\alpha-1)}\Bigr\rfloor=\Bigl\lfloor n\cdot(1{+}c)-\tfrac{1}{2}(2{-}\sqrt{2})\Bigr\rfloor$
of slope $\alpha/(\alpha-1)=1+c\approx 1.707$ ([4, Corollary 2.4]). The wall column $A_{m,0}$ is the non-homogeneous Beatty sequence
$A_{m,0}=\Bigl\lfloor\frac{m\alpha}{\alpha+1}\Bigr\rfloor=\Bigl\lfloor\frac{m}{\sqrt{2}}\Bigr\rfloor$ (24)
of slope $\alpha/(\alpha+1)=1/\sqrt{2}=c$ ([4, Corollary 2.5], and for $d=2$ this is A049472). Our canonical Beatty solution
$a_{\mathrm{B}}(n)=\bigl\lfloor(n{+}1)/\!\sqrt{2}\bigr\rfloor$
is a shifted version of the wall column (24), not of the first column. The difference word of $a_{\mathrm{B}}$ is Sturmian of slope $c$, and its multiplicity structure gives a Rayleigh–Fraenkel complementary partition of $\mathbb{Z}_{\geq 1}$ with slopes $1+\sqrt{2}$ and $(2+\sqrt{2})/2$ (see [5]).
Endpoint and defect interpretation
The endpoints of the interval $I$, their failure mechanisms, and the boundary defect set all read naturally in this framework. The left endpoint $d=\sqrt{2}/2$ is the unique shift for which the strong identity $a(f(n))=n$ holds, producing exactly the wall column (24) (shifted). The right endpoint $d=2(\sqrt{2}-1)=2/(1+\sqrt{2})$ is the maximal shift for which the addressing defect of the first nesting is still absorbed by the repeat structure of the underlying Sturmian difference word. Both endpoints lie in $\mathbb{Q}(\sqrt{2})\cap[0,1)$ with small Pell denominators. The failure of (5) at each endpoint is produced by orbit density near a single critical address $\theta=c$, as in Lemmas 17 and 18. The Pell-convergent identities
$cq_{k}=p_{k}+(-1)^{k}\varepsilon_{k},\qquad 2cp_{k}=q_{k}-2c(-1)^{k}\varepsilon_{k},$
with $\varepsilon_{k}=|cq_{k}-p_{k}|\to 0$, provide explicit Ostrowski witnesses of the dense orbit approaching $c$. The defect set at the upper endpoint, studied in Section 6, is the return-time set of the irrational rotation $\theta\mapsto\{\theta+c\}$ to the explicit interval [1c,1/2][1-c,1/2], and its gap alphabet $\{3,4,7\}$ consists of Pell-convergent denominators.
Generalisation to $r=s^{2}+1$
Remark 10 singled out the family $r=s^{2}+1$, $s\geq 1$, as the values of $r$ for which $N(r)=0$. This is exactly the family for which $c=1/\!\sqrt{r}$ has continued-fraction expansion $0;s,\overline{2s}]$ of minimal period one, and the corresponding Ostrowski numeration is Fokkink’s $m$-numeration with $m=2s$. The Pell tower ($s=1$) is the simplest case. Numerical exploration indicates that for every $s\geq 1$ the triple-nested identity admits a non-trivial interval of shifts. The endpoints and the associated defect structures for $s\geq 2$ are left for future work.
4.6 The triple-nested identity for $r=3$
The same mechanism extends to $r=3$. The three sliding terms produce a richer regime decomposition, and the repeat structure of $n$ and of n1n-1 no longer coincide, but the underlying logic (local regimes plus orbit density) gives a full if and only if.
Fix $r=3$, $c=1/\!\sqrt{3}$, and $a(n)=\lfloor cn+d\rfloor$. Write f(n)=a(n)+a(n1)+a(n2)f(n)=a(n)+a(n{-}1)+a(n{-}2) and $\Theta(n)=\theta_{n}+\theta_{n-1}+\theta_{n-2}$ with $\theta_{n}=\{cn+d\}$. Since $3c^{2}=1$,
$cf(n)+d=n-1+D-c\Theta(n),\qquad D=d(\sqrt{3}+1),$ (25)
and $a(f(n))=n-1+\lfloor D-c\Theta(n)\rfloor$. Setting $a(f(n))=n+m(n)$, the shift is
$m(n)=\lfloor D-c\Theta(n)\rfloor-1.$ (26)
As for $r=2$, the triple-nested identity $a(a(f(n)))=a(n)$ implies the two basic necessary conditions
  • (N+)
    $m(n)\geq 1\ \Rightarrow\ a(n+1)=a(n)$ ($n$ is a repeat),
  • (N-)
    $m(n)\leq-1\ \Rightarrow\ a(n-1)=a(n)$ (n1n-1 is a repeat),
but they are jointly equivalent to the identity only under the a priori bound $|m(n)|\leq 1$. We show below that this bound holds uniformly on $I_{3}$, so that on $I_{3}$ the necessity and the sufficiency parts can be read off from (N+) and (N-) together.